OpenStudy (anonymous):

Please Help! Find the solution set: (x^2-x)/(x^2-5x+6) >= 0

OpenStudy (anonymous):

?

OpenStudy (math&ing001):

f(x)=x^2-x=x(x-1) and g(x)=x^2-5x+6 Find the intervals in which f(x) and g(x) are both positive or both negative(*) and remove points where g(x)=0. That's your solution set. (*) because ++=+ and --=+

OpenStudy (anonymous):

so \[(-\infty, -1] \cup [2, +\infty) ???\]